OpenStudy (anonymous):

give me example problems about: deriving hyperbolic functions (please post it below)

OpenStudy (anonymous):

what is the value of \[\cosh^2 x-\sinh^2 x\]is a good example for workin on this concept

OpenStudy (anonymous):

it is equivalent to 1...

OpenStudy (anonymous):

i mean..... give me like a quiz about deriving hyperbolic functions...:)

OpenStudy (anonymous):

can u give me a simple one so i can give u similar one

hartnn (hartnn):

try to find sinh(A+B) = ? or is it too simple for you ?

OpenStudy (anonymous):

like: find y' of y = cosh (sin x)

hartnn (hartnn):

differentiate, \(\huge \frac{cosh\quad lnx+sinh\quad lnx}{cosh\quad lnx-sinh\quad lnx}\)

OpenStudy (anonymous):

wait

hartnn (hartnn):

waiting......

OpenStudy (anonymous):

not sure... \[\frac{ [\cosh(lnx)-\sinh(lnx)][\frac{ \sinh(lnx) }{ x }+\frac{ \cosh(lnx) }{ x }]-[\cosh(lnx)+\sinh(lnx)][\frac{ \sinh(lnx) }{ x }-\frac{ \cosh(lnx) }{ x }]}{ [\cosh(lnx)-\sinh(lnx)]^2 }\]

hartnn (hartnn):

lol,. the answer is 2x :P

hartnn (hartnn):

don't differentiate as it is, first simplify as much as u can.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

simplify? how?

hartnn (hartnn):

use cosh y as 1/2(e^y+e^-y) then use log property

OpenStudy (anonymous):

thanks!! more samples please?

hartnn (hartnn):

but could u do it ??

OpenStudy (anonymous):

f(x)= \[\frac{ e ^{lnx} }{ e ^{-lnx} }\] is that right?

hartnn (hartnn):

yup, what is e^{ln x} = ??

OpenStudy (anonymous):

? instead of x or y i used ln x <----given confused... what is your complete solution?

hartnn (hartnn):

e^{ln x} = x e^-{ln x} =1/ x x/(1/x) = x^2 d/dx(x^2) = 2x as simple as it is.

hartnn (hartnn):

next problem, differentiate tanh^-1 (sin x)

hartnn (hartnn):

or prove d/dx(tanh^-1 (sin x))=sec x

OpenStudy (anonymous):

we haven't study about inverse of hyperbolic func..

hartnn (hartnn):

integration ?

OpenStudy (anonymous):

just pre-cal...not yet on integration

hartnn (hartnn):

differentiate cosh^3 2x

hartnn (hartnn):

and then sinh (e^x)

OpenStudy (anonymous):

y'= 6 cosh^2 (2x) sinh (2x) y' =e^(x) [cosh (e^x)] correct?

hartnn (hartnn):

yup. diff. sqrt(coth 4x)

OpenStudy (anonymous):

y'=-4csch^2(4x)

hartnn (hartnn):

? no it was \(\sqrt{coth4x}\)

OpenStudy (anonymous):

lol..sorry y'=\[\frac{ -4csch^2(4x) }{ 2\sqrt{\coth(4x)} }\] ?

hartnn (hartnn):

yup. thats correct. next : differentiate cosh 5x sinh 3x

OpenStudy (anonymous):

\[y'= 5[\sinh(5x)\sinh(3x)]+3[\cosh(5x)\cosh(3x)]\] ?

hartnn (hartnn):

yup, now make one question for yourself and solve it.

OpenStudy (anonymous):

thanks a lot!! :))

hartnn (hartnn):

welcome :)